Determine the force of gravitational attraction between the earth (m=5.98×10²⁴kg) and a 45 kg physics student if the student is in a private jet together with his family at 6750 feet above earth's surface. This would place the student a distance of 7.39×10⁶m from earth's center

1
Expert's answer
2022-01-10T09:33:58-0500
"F=\dfrac{GM_Em}{R_E^2},""F=\dfrac{6.67\times10^{-11}\ \dfrac{N\times m^2}{kg^2}\times5.98\times10^{24}\ kg\times45\ kg}{(7.39\times10^6\ m)^2},""F=329\ N."